# SplatSuRe: Selective Super-Resolution for Multi-view Consistent 3D Gaussian Splatting

### v1

@@ A.4 Per-Scene Metrics
3DGS (LR) [kerbl3Dgaussians] 22.14 17.13 18.35 18.23 22.49 22.97 20.47 19.85 19.53 21.72 18.99 19.36 15.09 18.80 22.47 17.38 20.00 16.30 17.44
StableSR [wang2024exploiting] 23.62 20.52 19.68 22.20 25.23 25.18 21.65 23.24 21.75 24.72 20.24 22.44 22.54 20.55 25.97 22.86 22.02 22.72 19.76
Mip-Splatting [Yu2024MipSplatting] 23.46 21.22 18.24 23.15 27.11 26.85 21.29 24.13 22.26 26.16 20.31 22.99 24.09 21.33 26.62 23.94 21.99 23.48 20.28
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 24.23 21.26 19.69 23.44 26.98 26.83 21.74 24.12 22.74 26.17 20.49 22.98 23.82 21.28 26.55 23.87 22.54 23.72 20.63
Ours + StableSR [wang2024exploiting] 24.66 21.61 20.24 23.82 27.38 27.33 22.82 25.12 22.99 26.63 21.39 23.34 24.00 21.79 27.34 24.11 23.18 23.94 20.77
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
@@
3DGS (LR) [kerbl3Dgaussians] 26.29 27.15 18.50 22.75 23.27 17.60 18.64 19.55 25.93 20.34 19.44
StableSR [wang2024exploiting] 26.60 27.71 22.87 26.48 26.44 19.92 23.80 24.80 28.32 24.24 21.64
Mip-Splatting [Yu2024MipSplatting] 27.54 29.33 24.28 30.65 28.28 21.05 25.63 29.48 30.69 26.20 22.08
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 27.38 29.08 24.05 29.81 28.11 20.91 25.20 28.26 30.42 24.58 21.97
Ours + StableSR [wang2024exploiting] 28.46 29.57 24.04 30.54 28.39 20.55 25.31 29.56 31.05 25.61 22.02
Table 8: LPIPS $\downarrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]
@@
3DGS (LR) [kerbl3Dgaussians] 0.301 0.411 0.380 0.337 0.309 0.317 0.303 0.371 0.376 0.302 0.335 0.385 0.408 0.346 0.352 0.345 0.343 0.322 0.400
StableSR [wang2024exploiting] 0.247 0.357 0.364 0.283 0.275 0.295 0.283 0.333 0.294 0.266 0.302 0.340 0.289 0.305 0.316 0.298 0.306 0.234 0.317
Mip-Splatting [Yu2024MipSplatting] 0.275 0.331 0.408 0.288 0.260 0.293 0.296 0.316 0.323 0.250 0.325 0.325 0.281 0.298 0.316 0.284 0.313 0.237 0.345
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.242 0.328 0.367 0.266 0.251 0.277 0.276 0.311 0.292 0.242 0.299 0.321 0.271 0.289 0.304 0.274 0.294 0.221 0.315
Ours + StableSR [wang2024exploiting] 0.248 0.310 0.355 0.246 0.228 0.262 0.259 0.284 0.285 0.225 0.290 0.313 0.257 0.281 0.292 0.253 0.272 0.211 0.300
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
@@
3DGS (LR) [kerbl3Dgaussians] 0.332 0.339 0.424 0.328 0.306 0.477 0.400 0.312 0.293 0.436 0.482
StableSR [wang2024exploiting] 0.315 0.335 0.407 0.256 0.258 0.461 0.356 0.301 0.267 0.394 0.460
Mip-Splatting [Yu2024MipSplatting] 0.334 0.321 0.318 0.212 0.242 0.397 0.292 0.205 0.245 0.314 0.403
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.311 0.323 0.349 0.232 0.239 0.420 0.315 0.249 0.250 0.368 0.428
Ours + StableSR [wang2024exploiting] 0.300 0.312 0.368 0.238 0.241 0.454 0.328 0.202 0.253 0.371 0.452
Table 9: FID $\downarrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]
@@
3DGS (LR) [kerbl3Dgaussians] 86.75 85.58 83.25 73.89 40.93 40.48 84.46 82.51 99.19 53.16 81.79 29.93 113.45 43.80 29.64 35.69 58.24 148.48 88.82
StableSR [wang2024exploiting] 68.24 114.03 83.65 49.37 38.50 48.11 72.12 66.81 64.95 40.68 78.22 25.98 68.88 35.61 31.14 32.99 54.69 97.74 54.73
Mip-Splatting [Yu2024MipSplatting] 78.68 52.74 138.29 40.43 24.78 36.49 67.71 59.20 69.94 25.38 82.25 20.50 38.03 30.99 20.36 20.20 48.50 81.54 60.69
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 58.94 79.04 97.00 37.35 26.97 36.74 66.75 54.86 59.58 27.03 75.46 20.10 43.93 29.88 23.64 22.37 45.17 77.02 51.28
Ours + StableSR [wang2024exploiting] 60.48 46.33 70.53 27.26 20.96 27.20 57.93 35.69 54.34 19.79 63.87 16.42 28.28 22.83 16.72 14.82 32.61 58.61 41.93
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
@@
3DGS (LR) [kerbl3Dgaussians] 60.62 72.04 30.50 72.93 79.04 77.01 47.98 59.89 41.92 89.83 42.08
StableSR [wang2024exploiting] 51.64 67.41 23.55 54.42 64.89 73.04 19.38 49.64 37.09 105.27 40.16
Mip-Splatting [Yu2024MipSplatting] 53.90 50.26 10.21 20.19 36.29 22.76 9.05 12.04 17.57 37.03 20.21
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 47.05 51.10 11.37 26.41 41.87 46.01 10.46 15.43 21.62 42.73 26.52
Ours + StableSR [wang2024exploiting] 43.19 45.09 11.79 22.40 38.03 58.56 10.63 8.47 17.69 50.63 29.83
Table 10: CMMD $\downarrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]
@@
3DGS (LR) [kerbl3Dgaussians] 1.770 2.234 2.361 1.602 0.718 1.699 1.540 1.925 1.616 1.202 2.522 1.897 3.038 2.153 2.758 2.140 1.705 3.211 2.165
StableSR [wang2024exploiting] 1.093 1.032 1.491 1.073 0.791 1.068 1.092 1.139 0.987 0.907 1.597 1.464 0.647 1.330 1.401 1.143 1.163 0.779 1.142
Mip-Splatting [Yu2024MipSplatting] 1.246 1.046 1.624 0.787 0.633 1.140 0.994 1.151 1.024 0.579 1.853 1.243 0.863 1.384 1.286 1.080 1.029 1.105 1.528
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 1.066 0.931 1.582 0.787 0.516 0.942 0.978 0.998 0.943 0.523 1.795 1.261 0.557 1.383 1.408 1.140 1.049 0.658 1.400
Ours + StableSR [wang2024exploiting] 0.993 1.121 1.568 0.643 0.351 1.134 0.901 1.025 0.951 0.443 1.699 1.330 0.521 1.529 1.355 1.061 0.944 0.897 1.286
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
@@
3DGS (LR) [kerbl3Dgaussians] 1.003 0.733 0.392 1.011 0.319 1.354 0.242 0.346 0.261 0.926 0.956
StableSR [wang2024exploiting] 0.807 0.714 0.621 0.934 0.440 1.409 0.310 0.441 0.578 0.767 1.023
Mip-Splatting [Yu2024MipSplatting] 0.828 0.552 0.099 0.346 0.195 0.215 0.036 0.191 0.187 0.125 0.257
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.715 0.546 0.138 0.417 0.279 0.782 0.081 0.228 0.378 0.261 0.486
Ours + StableSR [wang2024exploiting] 0.558 0.434 0.159 0.266 0.139 1.108 0.065 0.119 0.184 0.406 0.610
Table 11: DreamSim $\downarrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]
@@
3DGS (LR) [kerbl3Dgaussians] 0.0969 0.0884 0.1548 0.0881 0.0399 0.0549 0.1154 0.0806 0.1074 0.0552 0.1156 0.0595 0.1215 0.0747 0.0710 0.0620 0.0795 0.1142 0.1207
StableSR [wang2024exploiting] 0.0634 0.0703 0.1497 0.0438 0.0443 0.0613 0.0909 0.0759 0.0566 0.0591 0.0955 0.0516 0.0565 0.0652 0.0555 0.0590 0.0558 0.0505 0.0627
Mip-Splatting [Yu2024MipSplatting] 0.0739 0.0465 0.2459 0.0312 0.0216 0.0392 0.0887 0.0541 0.0592 0.0266 0.1001 0.0392 0.0278 0.0479 0.0341 0.0293 0.0541 0.0438 0.0703
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.0500 0.0511 0.1603 0.0271 0.0264 0.0399 0.0792 0.0564 0.0466 0.0369 0.0936 0.0416 0.0359 0.0504 0.0402 0.0402 0.0459 0.0390 0.0567
Ours + StableSR [wang2024exploiting] 0.0490 0.0381 0.1272 0.0202 0.0187 0.0314 0.0625 0.0369 0.0422 0.0239 0.0720 0.0315 0.0229 0.0410 0.0309 0.0268 0.0331 0.0279 0.0486
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
@@
3DGS (LR) [kerbl3Dgaussians] 0.0613 0.0534 0.0624 0.0659 0.0453 0.0562 0.0452 0.0454 0.0322 0.0586 0.0558
StableSR [wang2024exploiting] 0.0499 0.0510 0.0348 0.0320 0.0377 0.0598 0.0170 0.0240 0.0345 0.0536 0.0553
Mip-Splatting [Yu2024MipSplatting] 0.0451 0.0345 0.0126 0.0122 0.0166 0.0124 0.0070 0.0046 0.0163 0.0117 0.0252
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.0439 0.0378 0.0150 0.0160 0.0226 0.0331 0.0079 0.0066 0.0212 0.0178 0.0345
Ours + StableSR [wang2024exploiting] 0.0332 0.0327 0.0151 0.0123 0.0177 0.0380 0.0071 0.0037 0.0144 0.0168 0.0364
Table 12: MUSIQ $\uparrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]
@@
3DGS (LR) [kerbl3Dgaussians] 55.443 64.965 51.161 56.309 54.610 60.048 48.785 54.510 55.349 53.631 57.565 59.116 65.987 60.573 59.605 61.675 57.991 57.572 62.845
StableSR [wang2024exploiting] 58.573 58.466 49.326 56.501 57.051 60.822 46.028 53.817 55.209 54.985 54.748 60.520 62.883 60.841 55.264 60.637 55.297 54.926 62.311
Mip-Splatting [Yu2024MipSplatting] 51.773 48.745 42.653 50.651 46.800 49.040 33.568 41.815 45.084 47.621 46.236 51.797 49.461 51.300 40.700 47.208 47.433 39.068 53.893
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 57.276 57.690 48.339 54.321 54.969 58.305 44.228 52.796 54.435 53.662 52.762 58.732 61.687 58.721 54.220 57.781 54.378 53.369 61.291
Ours + StableSR [wang2024exploiting] 58.504 60.078 50.762 59.501 56.915 61.660 47.882 58.036 57.351 55.500 55.978 60.757 65.967 60.214 57.281 62.616 57.574 58.244 63.491
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
@@
3DGS (LR) [kerbl3Dgaussians] 47.916 49.425 54.239 56.108 55.829 58.968 58.961 63.552 50.931 52.085 41.783
StableSR [wang2024exploiting] 52.919 53.324 55.767 59.725 57.965 52.194 54.576 59.952 58.135 44.337 38.276
Mip-Splatting [Yu2024MipSplatting] 38.800 44.105 38.243 44.656 42.445 35.728 44.342 50.600 41.878 35.866 35.112
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 50.849 52.035 54.818 56.237 55.391 53.542 54.318 58.321 55.637 44.039 37.260
Ours + StableSR [wang2024exploiting] 50.594 55.444 57.923 58.501 57.752 51.206 56.236 64.952 55.154 49.551 38.014
Table 13: NIQE $\downarrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]
@@
3DGS (LR) [kerbl3Dgaussians] 4.718 2.628 4.293 2.791 3.884 3.944 4.097 2.744 3.385 3.625 4.081 2.873 2.605 2.994 4.269 2.784 3.169 3.011 2.926
StableSR [wang2024exploiting] 6.531 4.545 5.820 4.335 5.239 5.229 5.415 4.991 4.830 5.022 4.844 4.690 4.214 4.363 5.611 4.361 5.005 4.564 4.356
Mip-Splatting [Yu2024MipSplatting] 5.482 4.693 6.567 4.013 5.614 5.302 5.482 5.086 4.569 4.955 5.039 4.632 5.077 4.449 5.982 5.022 4.653 5.027 4.179
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 6.128 4.187 5.629 3.978 4.904 4.988 5.025 4.537 4.562 4.716 4.573 4.323 4.026 4.137 5.234 4.096 4.644 4.280 4.050
Ours + StableSR [wang2024exploiting] 5.330 3.440 4.720 3.318 3.997 4.657 4.309 3.644 4.029 3.765 4.167 3.340 3.435 3.498 4.559 3.416 3.785 3.632 3.582
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
@@
3DGS (LR) [kerbl3Dgaussians] 4.760 5.329 2.808 3.964 3.535 2.915 2.741 2.986 4.210 3.137 2.744
StableSR [wang2024exploiting] 5.231 6.225 4.333 5.452 5.533 4.938 5.516 5.633 6.013 4.919 5.105
Mip-Splatting [Yu2024MipSplatting] 6.097 6.426 5.665 6.458 5.730 5.965 6.155 5.600 5.885 5.623 5.462
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 5.208 6.188 4.131 5.228 5.214 4.521 5.074 5.546 5.840 4.846 4.381
Ours + StableSR [wang2024exploiting] 5.054 5.768 3.378 4.309 3.828 3.825 3.949 3.828 4.531 4.052 3.702

### v2

@@ A.5 Per-Scene Metrics
3DGS (LR) [14] 22.14 17.13 18.35 18.23 22.49 22.97 20.47 19.85 19.53 21.72 18.99 19.36 15.09 18.80 22.47 17.38 20.00 16.30 17.44
3DGS [14] + StableSR [27] 23.62 20.52 19.68 22.20 25.23 25.18 21.65 23.24 21.75 24.72 20.24 22.44 22.54 20.55 25.97 22.86 22.02 22.72 19.76
Mip-Splatting [33] 23.46 21.22 18.24 23.15 27.11 26.85 21.29 24.13 22.26 26.16 20.31 22.99 24.09 21.33 26.62 23.94 21.99 23.48 20.28
SRGS [6] + StableSR [27] 24.23 21.26 19.69 23.44 26.98 26.83 21.74 24.12 22.74 26.17 20.49 22.98 23.82 21.28 26.55 23.87 22.54 23.72 20.63
Ours + StableSR [27] 24.66 21.61 20.24 23.82 27.38 27.33 22.82 25.12 22.99 26.63 21.39 23.34 24.00 21.79 27.34 24.11 23.18 23.94 20.77
Method Deep Blending [8] Mip-NeRF 360 [1]
@@
3DGS (LR) [14] 26.29 27.15 18.50 22.75 23.27 17.60 18.64 19.55 25.93 20.34 19.44
3DGS [14] + StableSR [27] 26.60 27.71 22.87 26.48 26.44 19.92 23.80 24.80 28.32 24.24 21.64
Mip-Splatting [33] 27.54 29.33 24.28 30.65 28.28 21.05 25.63 29.48 30.69 26.20 22.08
SRGS [6] + StableSR [27] 27.38 29.08 24.05 29.81 28.11 20.91 25.20 28.26 30.42 24.58 21.97
Ours + StableSR [27] 28.46 29.57 24.04 30.54 28.39 20.55 25.31 29.56 31.05 25.61 22.02
Table 9: LPIPS $\downarrow$ on each scene in Tanks & Temples [15], Deep Blending [8], and Mip-NeRF 360 [1].
Method Tanks & Temples [15]
@@
3DGS (LR) [14] 0.301 0.411 0.380 0.337 0.309 0.317 0.303 0.371 0.376 0.302 0.335 0.385 0.408 0.346 0.352 0.345 0.343 0.322 0.400
3DGS [14] + StableSR [27] 0.247 0.357 0.364 0.283 0.275 0.295 0.283 0.333 0.294 0.266 0.302 0.340 0.289 0.305 0.316 0.298 0.306 0.234 0.317
Mip-Splatting [33] 0.275 0.331 0.408 0.288 0.260 0.293 0.296 0.316 0.323 0.250 0.325 0.325 0.281 0.298 0.316 0.284 0.313 0.237 0.345
SRGS [6] + StableSR [27] 0.242 0.328 0.367 0.266 0.251 0.277 0.276 0.311 0.292 0.242 0.299 0.321 0.271 0.289 0.304 0.274 0.294 0.221 0.315
Ours + StableSR [27] 0.248 0.310 0.355 0.246 0.228 0.262 0.259 0.284 0.285 0.225 0.290 0.313 0.257 0.281 0.292 0.253 0.272 0.211 0.300
Method Deep Blending [8] Mip-NeRF 360 [1]
@@
3DGS (LR) [14] 0.332 0.339 0.424 0.328 0.306 0.477 0.400 0.312 0.293 0.436 0.482
3DGS [14] + StableSR [27] 0.315 0.335 0.407 0.256 0.258 0.461 0.356 0.301 0.267 0.394 0.460
Mip-Splatting [33] 0.334 0.321 0.318 0.212 0.242 0.397 0.292 0.205 0.245 0.314 0.403
SRGS [6] + StableSR [27] 0.311 0.323 0.349 0.232 0.239 0.420 0.315 0.249 0.250 0.368 0.428
Ours + StableSR [27] 0.300 0.312 0.368 0.238 0.241 0.454 0.328 0.202 0.253 0.371 0.452
Table 10: FID $\downarrow$ on each scene in Tanks & Temples [15], Deep Blending [8], and Mip-NeRF 360 [1].
Method Tanks & Temples [15]
@@
3DGS (LR) [14] 86.75 85.58 83.25 73.89 40.93 40.48 84.46 82.51 99.19 53.16 81.79 29.93 113.45 43.80 29.64 35.69 58.24 148.48 88.82
3DGS [14] + StableSR [27] 68.24 114.03 83.65 49.37 38.50 48.11 72.12 66.81 64.95 40.68 78.22 25.98 68.88 35.61 31.14 32.99 54.69 97.74 54.73
Mip-Splatting [33] 78.68 52.74 138.29 40.43 24.78 36.49 67.71 59.20 69.94 25.38 82.25 20.50 38.03 30.99 20.36 20.20 48.50 81.54 60.69
SRGS [6] + StableSR [27] 58.94 79.04 97.00 37.35 26.97 36.74 66.75 54.86 59.58 27.03 75.46 20.10 43.93 29.88 23.64 22.37 45.17 77.02 51.28
Ours + StableSR [27] 60.48 46.33 70.53 27.26 20.96 27.20 57.93 35.69 54.34 19.79 63.87 16.42 28.28 22.83 16.72 14.82 32.61 58.61 41.93
Method Deep Blending [8] Mip-NeRF 360 [1]
@@
3DGS (LR) [14] 60.62 72.04 30.50 72.93 79.04 77.01 47.98 59.89 41.92 89.83 42.08
3DGS [14] + StableSR [27] 51.64 67.41 23.55 54.42 64.89 73.04 19.38 49.64 37.09 105.27 40.16
Mip-Splatting [33] 53.90 50.26 10.21 20.19 36.29 22.76 9.05 12.04 17.57 37.03 20.21
SRGS [6] + StableSR [27] 47.05 51.10 11.37 26.41 41.87 46.01 10.46 15.43 21.62 42.73 26.52
Ours + StableSR [27] 43.19 45.09 11.79 22.40 38.03 58.56 10.63 8.47 17.69 50.63 29.83
Table 11: CMMD $\downarrow$ on each scene in Tanks & Temples [15], Deep Blending [8], and Mip-NeRF 360 [1].
Method Tanks & Temples [15]
@@
3DGS (LR) [14] 1.770 2.234 2.361 1.602 0.718 1.699 1.540 1.925 1.616 1.202 2.522 1.897 3.038 2.153 2.758 2.140 1.705 3.211 2.165
3DGS [14] + StableSR [27] 1.093 1.032 1.491 1.073 0.791 1.068 1.092 1.139 0.987 0.907 1.597 1.464 0.647 1.330 1.401 1.143 1.163 0.779 1.142
Mip-Splatting [33] 1.246 1.046 1.624 0.787 0.633 1.140 0.994 1.151 1.024 0.579 1.853 1.243 0.863 1.384 1.286 1.080 1.029 1.105 1.528
SRGS [6] + StableSR [27] 1.066 0.931 1.582 0.787 0.516 0.942 0.978 0.998 0.943 0.523 1.795 1.261 0.557 1.383 1.408 1.140 1.049 0.658 1.400
Ours + StableSR [27] 0.993 1.121 1.568 0.643 0.351 1.134 0.901 1.025 0.951 0.443 1.699 1.330 0.521 1.529 1.355 1.061 0.944 0.897 1.286
Method Deep Blending [8] Mip-NeRF 360 [1]
@@
3DGS (LR) [14] 1.003 0.733 0.392 1.011 0.319 1.354 0.242 0.346 0.261 0.926 0.956
3DGS [14] + StableSR [27] 0.807 0.714 0.621 0.934 0.440 1.409 0.310 0.441 0.578 0.767 1.023
Mip-Splatting [33] 0.828 0.552 0.099 0.346 0.195 0.215 0.036 0.191 0.187 0.125 0.257
SRGS [6] + StableSR [27] 0.715 0.546 0.138 0.417 0.279 0.782 0.081 0.228 0.378 0.261 0.486
Ours + StableSR [27] 0.558 0.434 0.159 0.266 0.139 1.108 0.065 0.119 0.184 0.406 0.610
Table 12: DreamSim $\downarrow$ on each scene in Tanks & Temples [15], Deep Blending [8], and Mip-NeRF 360 [1].
Method Tanks & Temples [15]
@@
3DGS (LR) [14] 0.0969 0.0884 0.1548 0.0881 0.0399 0.0549 0.1154 0.0806 0.1074 0.0552 0.1156 0.0595 0.1215 0.0747 0.0710 0.0620 0.0795 0.1142 0.1207
3DGS [14] + StableSR [27] 0.0634 0.0703 0.1497 0.0438 0.0443 0.0613 0.0909 0.0759 0.0566 0.0591 0.0955 0.0516 0.0565 0.0652 0.0555 0.0590 0.0558 0.0505 0.0627
Mip-Splatting [33] 0.0739 0.0465 0.2459 0.0312 0.0216 0.0392 0.0887 0.0541 0.0592 0.0266 0.1001 0.0392 0.0278 0.0479 0.0341 0.0293 0.0541 0.0438 0.0703
SRGS [6] + StableSR [27] 0.0500 0.0511 0.1603 0.0271 0.0264 0.0399 0.0792 0.0564 0.0466 0.0369 0.0936 0.0416 0.0359 0.0504 0.0402 0.0402 0.0459 0.0390 0.0567
Ours + StableSR [27] 0.0490 0.0381 0.1272 0.0202 0.0187 0.0314 0.0625 0.0369 0.0422 0.0239 0.0720 0.0315 0.0229 0.0410 0.0309 0.0268 0.0331 0.0279 0.0486
Method Deep Blending [8] Mip-NeRF 360 [1]
@@
3DGS (LR) [14] 0.0613 0.0534 0.0624 0.0659 0.0453 0.0562 0.0452 0.0454 0.0322 0.0586 0.0558
3DGS [14] + StableSR [27] 0.0499 0.0510 0.0348 0.0320 0.0377 0.0598 0.0170 0.0240 0.0345 0.0536 0.0553
Mip-Splatting [33] 0.0451 0.0345 0.0126 0.0122 0.0166 0.0124 0.0070 0.0046 0.0163 0.0117 0.0252
SRGS [6] + StableSR [27] 0.0439 0.0378 0.0150 0.0160 0.0226 0.0331 0.0079 0.0066 0.0212 0.0178 0.0345
Ours + StableSR [27] 0.0332 0.0327 0.0151 0.0123 0.0177 0.0380 0.0071 0.0037 0.0144 0.0168 0.0364
Table 13: MUSIQ $\uparrow$ on each scene in Tanks & Temples [15], Deep Blending [8], and Mip-NeRF 360 [1].
Method Tanks & Temples [15]
@@
3DGS (LR) [14] 55.443 64.965 51.161 56.309 54.610 60.048 48.785 54.510 55.349 53.631 57.565 59.116 65.987 60.573 59.605 61.675 57.991 57.572 62.845
3DGS [14] + StableSR [27] 58.573 58.466 49.326 56.501 57.051 60.822 46.028 53.817 55.209 54.985 54.748 60.520 62.883 60.841 55.264 60.637 55.297 54.926 62.311
Mip-Splatting [33] 51.773 48.745 42.653 50.651 46.800 49.040 33.568 41.815 45.084 47.621 46.236 51.797 49.461 51.300 40.700 47.208 47.433 39.068 53.893
SRGS [6] + StableSR [27] 57.276 57.690 48.339 54.321 54.969 58.305 44.228 52.796 54.435 53.662 52.762 58.732 61.687 58.721 54.220 57.781 54.378 53.369 61.291
Ours + StableSR [27] 58.504 60.078 50.762 59.501 56.915 61.660 47.882 58.036 57.351 55.500 55.978 60.757 65.967 60.214 57.281 62.616 57.574 58.244 63.491
Method Deep Blending [8] Mip-NeRF 360 [1]
@@
3DGS (LR) [14] 47.916 49.425 54.239 56.108 55.829 58.968 58.961 63.552 50.931 52.085 41.783
3DGS [14] + StableSR [27] 52.919 53.324 55.767 59.725 57.965 52.194 54.576 59.952 58.135 44.337 38.276
Mip-Splatting [33] 38.800 44.105 38.243 44.656 42.445 35.728 44.342 50.600 41.878 35.866 35.112
SRGS [6] + StableSR [27] 50.849 52.035 54.818 56.237 55.391 53.542 54.318 58.321 55.637 44.039 37.260
Ours + StableSR [27] 50.594 55.444 57.923 58.501 57.752 51.206 56.236 64.952 55.154 49.551 38.014
Table 14: NIQE $\downarrow$ on each scene in Tanks & Temples [15], Deep Blending [8], and Mip-NeRF 360 [1].
Method Tanks & Temples [15]
@@
3DGS (LR) [14] 4.718 2.628 4.293 2.791 3.884 3.944 4.097 2.744 3.385 3.625 4.081 2.873 2.605 2.994 4.269 2.784 3.169 3.011 2.926
3DGS [14] + StableSR [27] 6.531 4.545 5.820 4.335 5.239 5.229 5.415 4.991 4.830 5.022 4.844 4.690 4.214 4.363 5.611 4.361 5.005 4.564 4.356
Mip-Splatting [33] 5.482 4.693 6.567 4.013 5.614 5.302 5.482 5.086 4.569 4.955 5.039 4.632 5.077 4.449 5.982 5.022 4.653 5.027 4.179
SRGS [6] + StableSR [27] 6.128 4.187 5.629 3.978 4.904 4.988 5.025 4.537 4.562 4.716 4.573 4.323 4.026 4.137 5.234 4.096 4.644 4.280 4.050
Ours + StableSR [27] 5.330 3.440 4.720 3.318 3.997 4.657 4.309 3.644 4.029 3.765 4.167 3.340 3.435 3.498 4.559 3.416 3.785 3.632 3.582
Method Deep Blending [8] Mip-NeRF 360 [1]
@@
3DGS (LR) [14] 4.760 5.329 2.808 3.964 3.535 2.915 2.741 2.986 4.210 3.137 2.744
3DGS [14] + StableSR [27] 5.231 6.225 4.333 5.452 5.533 4.938 5.516 5.633 6.013 4.919 5.105
Mip-Splatting [33] 6.097 6.426 5.665 6.458 5.730 5.965 6.155 5.600 5.885 5.623 5.462
SRGS [6] + StableSR [27] 5.208 6.188 4.131 5.228 5.214 4.521 5.074 5.546 5.840 4.846 4.381
Ours + StableSR [27] 5.054 5.768 3.378 4.309 3.828 3.825 3.949 3.828 4.531 4.052 3.702